# Review of Demographic Fairness in Face Recognition

### v2

@@ Abstract
Demographic fairness in face recognition (FR) has emerged as a critical area of research, given its impact on fairness, equity, and reliability across diverse applications. As FR technologies are increasingly deployed globally, disparities in performance across demographic groups– such as race, ethnicity, and gender– have garnered significant attention. These biases not only compromise the credibility of FR systems but also raise ethical concerns, especially when these technologies are employed in sensitive domains. This review consolidates extensive research efforts providing a comprehensive overview of the multifaceted aspects of demographic fairness in FR.
We systematically examine the primary causes, datasets, assessment metrics, and mitigation approaches associated with demographic disparities in FR. By categorizing key contributions in these areas, this work provides a structured approach to understanding and addressing the complexity of this issue. Finally, we highlight current advancements and identify emerging challenges that need further investigation. This article aims to provide researchers with a unified perspective on the state-of-the-art while emphasizing the critical need for equitable and trustworthy FR systems.
@@ I Introduction
Demographic fairness in face recognition (FR) systems has emerged as a critical challenge in the deployment of biometric technologies for real-world applications [1, 2, 3, 4]. Disparities in these systems often leads to disparities in performance across demographic groups- such as variations in recognition accuracy- based on race, gender, and age [4, 5, 6]. Such disparities, also commonly referred to as biases, can have far-reaching consequences, especially in critical applications like border crossing, law enforcement [7, 8], security [9], and hiring processes [10, 11, 12], where fairness and accuracy are paramount.
A conventional FR system evaluates the similarity between two facial images by computing a matching score based on their extracted features to determine whether they correspond to the same individual or not. The performance of such systems is typically examined by analyzing the distributions of matching scores for image pairs of the same identity (referred to as mated pairs) and those of different identities (non-mated pairs). Scores are thresholded into binarized decisions: match or no-match. Fig. 1 illustrates the issue of demographic bias based on race, gender, and age. Ideally, a fair FR model should exhibit equitable performance across all demographic groups (i.e., similar error rates or comparable score distributions). The disparity between groups contributes to the unfair or biased FR. (The formal definitions for error rates and disparities provided in Sec. IV.) As depicted by various choices of scoring thresholds, accuracy differentials exist across demographic splits and can manifest at various thresholds. This emphasizes that demographic disparities are not tied solely to false match or false non-match errors but may be present throughout the score space. This issue is further compounded by the growing reliance on FR technologies, making it imperative to identify, evaluate, and mitigate sources of accuracy differentials effectively.
@@
Figure 1: Illustration of differential performance across demographic groups in FR with reference to different demographic factors: (a) race or ethnicity, (b) gender, and (c) age. Score distributions for mated and non-mated pairs are shown for multiple demographic groups, along with decision thresholds (dotted lines) corresponding to different operating points. This figure is intended to demonstrate that accuracy differentials may persist across a range of thresholds and are not restricted to a particular type of error (e.g., false match or false non-match).
@@
Due to its severity and the widespread range of FR applications, demographic fairness has emerged as a crucial area of research, drawing significant attention from both the biometrics and computer vision communities [3, 4, 13, 14]. This issue has been formally incorporated into the evaluation frameworks of prominent initiatives, such as the National Institute of Standards and Technology’s (NIST) Face Recognition Vendor Tests (FRVT), which have included demographic effects in their reports since 2019 [15, 16]. Other initiatives, including the Maryland Test Facility (MdTF) and the European Association for Biometrics (EAB), have also contributed to this discourse, though at a smaller scale and scope compared to NIST. The MdTF, supported by the United States Department of Homeland Security (DHS), has conducted biometric technology rallies to evaluate demographic disparities in FR systems. In Europe, organizations such as the EAB have hosted dedicated events on demographic fairness in biometric systems, underlining the global importance of this topic. This research area is often positioned within the broader context of fair and trustworthy biometrics, and has been receiving substantial attention– in the form of papers, workshops, or special sessions– from leading conferences such as IEEE/CVF CVPR, WACV, IEEE FG, and ICPR, and from reputable journals including IEEE Transactions on Information Forensics and Security (TIFS), IEEE Transactions on Biometrics, Behavior, and Identity Science (TBIOM) [17] and IEEE Signal Processing Magazine [18]. Additionally, standards organizations have recognized the need for systematic approaches. For example, the recently published ISO/IEC 19795-10 guidelines quantify demographic differentials in biometric systems, and emphasize the need for addressing fairness in FR technologies [19].
@@
Scope: Fairness and bias in machine learning are expansive topics, and their application in biometrics has drawn significant attention in recent years. Several comprehensive reviews have addressed fairness and bias in machine learning broadly [5, 20], while others focus specifically on biometrics, offering insights into various modalities such as face, fingerprint, and vein, alongside applications beyond recognition, including region of interest (ROI) detection, quality assessment, and presentation attack detection[3, 4, 13, 21]. However, as face remains the most commonly used biometric trait in operational scenarios, a substantial portion of research on demographic bias has concentrated on this modality.
@@
While previous reviews offer broad perspectives, the extensive literature and emerging challenges specific to demographic fairness in FR necessitate a dedicated review. In this article, we provide a consolidated discussion of recent advancements in the field, addressing the causes of demographic bias, available datasets for research, evaluation metrics for assessment of fairness, and recent mitigation techniques. Furthermore, we explore ongoing challenges that persist in addressing demographic disparities, particularly in the light of novel use cases and emerging FR applications. Although our primary focus is on race and ethnicity, as these are dominant areas of research, we also include gender-related studies within the broader context. Age-related studies, given their distinct nature and established body of research, are referenced only where directly relevant.
Naming Conventions: In this review, we use the term demographic bias to refer to accuracy differentials observed across demographic groups (e.g., race, gender, age). We adopt this terminology to remain consistent with much of the recent literature in FR and biometrics community. It may be noted that this term is not intended to imply intentional discrimination or moral judgment; rather, it reflects measurable differences in algorithmic performance across groupings that are often of regulatory and ethical interest.
@@
In this review, we adopt the terminologies used in the referenced works while acknowledging the minor differences in naming conventions. Terms such as race and ethnicity are often used interchangeably, although they represent distinct concepts. For clarity, we retain the terms employed by the original studies. Similarly, the names of ethnic groups vary across the literature: for instance, some works use terms like Black and White, while others prefer African (or African-American) and Caucasian. Additionally, the South Asian group is sometimes referred to as Indian, whereas Asians often refer to East Asians. To maintain consistency and respect the source material, we adhere to the original terminologies in this review. Readers are encouraged to refer to the cited works for precise definitions and context. While fairness in related topics such as face detection, image quality, expression recognition, and attribute estimation is of interest to the research community, this review exclusively focuses on demographic fairness in FR.
@@
Contributions: This review constitutes the first comprehensive work dedicated to exploring demographic fairness in FR, offering a unified and holistic perspective to researchers in the field. We systematically analyze and organize key aspects, including the causes of demographic disparities, available datasets, assessment metrics, and mitigation techniques, providing a structured framework for understanding these areas. Finally, we identify emerging challenges and unresolved questions, inviting further research and innovation to advance fairness in FR systems.
@@
The structure of this paper is as follows: Section II explores the causes of bias, analyzing factors such as distribution of demographic groups in datasets, skin-tone, image quality, and soft attributes. We provide an overview of datasets commonly used for fairness-related research in Section III, highlighting their demographic attributes and suitability for specific tasks. Section IV reviews existing metrics for fairness evaluation, discussing their strengths and limitations. Section V outlines recent bias mitigation strategies across different stages of the FR pipeline. We discuss open challenges and future research directions in Section VI, and conclude the review in Section VII.
II Causes of Demographic Differentials
@@
In this section, we consolidate findings from existing works related to causes of accuracy differentials in FR. Considering the wide range of research in this area, we have grouped the causes into categories for clarity and systematic review. These categories encompass factors such as imbalances in training datasets, variability in skin-tones, algorithmic sensitivity, image quality, related covariates, combined or intersectional factors, and soft attributes. Though this categorization simplifies the organization, it is important to note that many studies attribute accuracy differentials to multiple, overlapping factors, making strict classification difficult. We have categorized works based on their primary focus or findings, but some studies have been referenced across multiple categories to reflect their broader relevance. This approach also highlights that causes of demographic differentials or bias are inherently multifaceted and interconnected, requiring a thorough understanding to analyze and mitigate these disparities. A quick summary of various works discussed in this section can be found in Table I.
II-A Training Datasets
@@
Early studies in this area speculated that the issue of demographic bias in FR systems often stems from imbalanced or unrepresentative datasets, significantly influencing both training and evaluation outcomes. Research by Krishnapriya et al. [22] demonstrated how demographic groups, such as African-American cohorts, exhibit higher false match rates (FMR), while Caucasian cohorts face higher false non-match rates (FNMR), highlighting the interplay between race and matching thresholds. The Face Recognition Vendor Test (FRVT) conducted by NIST [15] substantiated these findings, reporting increased false positive identifications in women, children, and the elderly, alongside higher false negatives in under-represented racial groups, emphasizing the intricate interactions between dataset characteristics and demographic attributes.
@@
In a pre-deep learning era, studies like Klare et al.  [23] examined the impact of demographic balance in training datasets on FR performance. While their results indicated that underrepresented groups– such as females and Black individuals– benefited from increased representation, they also found that balancing the dataset alone did not fully eliminate disparities in recognition accuracy at test time. To further mitigate this, they proposed training exclusive cohort-specific models and using dynamic model selection during inference, highlighting that data balance might often be complemented with architectural or procedural adjustments to meaningfully reduce bias. Cavazos et al. [24] identified how dataset complexity and identification thresholds contribute to racial bias, such as the need for higher thresholds for East-Asian faces to achieve comparable false acceptance rates (FARs). Gwilliam et al. [25] challenged the prevailing assumptions about the necessity of balanced datasets by demonstrating that skewed distributions favoring African faces reduced racial differentials more effectively than balanced datasets. In [26, 27], Albiero et al. demonstrated that balancing training data for both genders (male and female) did not help reducing the corresponding differential in FR accuracy. Wu and Bowyer [28] expanded this discussion, emphasizing that mere balance in identities or number of images is insufficient to address disparities, highlighting additional factors like brightness and head pose during dataset assembly.
@@
Other works delved into specific aspects of demographic balance in datasets. Wang et al.  [29] observed that even race-balanced datasets failed to eliminate accuracy differentials, hypothesizing that certain ethnicities are inherently more challenging to recognize. Kolla and Savadamuthu [30] highlighted the influence of facial quality and racial feature gradations on model fairness. Focusing on inter-sectional bias, Muthukumar et al.  [31] identified structural facial features as significant contributors, particularly for dark-skinned females, over attributes like skin-tone or hair length. Cook et al. [32] further analyzed the role of image acquisition conditions, noting how factors such as skin reflectance and environmental conditions disproportionately affect darker-skinned individuals, thus advocating for standardized acquisition protocols to mitigate bias.
@@
II-B Variability in Skin-tone

### v3

@@ Abstract
The issue of difference in face recognition (FR) performance across demographic groups has emerged as a critical area of research, given its impact on fairness, equity, and reliability across diverse applications. As FR technologies are increasingly deployed globally, disparities in performance across demographic groups– such as race, ethnicity, and gender– have garnered significant attention. These differences or biases not only compromise the credibility of FR systems but also raise ethical concerns, especially when these technologies are employed in sensitive domains. This review consolidates extensive research efforts providing a comprehensive overview of the multifaceted aspects of demographic fairness in FR.
We systematically examine the primary causes, datasets, assessment metrics, and mitigation approaches associated with performance differences in FR across demographic groups. By categorizing key contributions in these areas, this work provides a structured approach to understanding and addressing the complexity of this issue. Finally, we highlight current advancements and identify emerging challenges that need further investigation. This article aims to provide researchers with a unified perspective on the state-of-the-art while emphasizing the critical need for equitable and trustworthy FR systems.
Index Terms:
Demographic Fairness, Bias, Face Recognition, Biometrics, Differential Performance, Trustworthy AI
@@ I Introduction
Differences in the performance of face recognition (FR) system across demographic groups, commonly described under the umbrella of demographic fairness, have emerged as a critical challenge in the deployment of biometric technologies for real-world applications [1, 2, 3, 4]. Several real-world incidents underscore the societal risks associated with such disparities. For example, a Black individual was wrongly arrested in Detroit due to a false FR match [5]; a US civil liberties test reported misidentification of a handful US lawmakers as criminals by major FR service [6]; and recently, researchers exposed racial bias in airport FR systems used during passenger boarding [7, 8]. These systems frequently exhibit unequal recognition performance across race, gender, and age groups [4, 9, 10]. Such disparities, often labelled as biases, can have far-reaching consequences, especially in critical applications like border crossing, law enforcement [11, 12], security [13], and hiring processes [14, 15, 16], where both fairness and accuracy are paramount.
@@
Figure 1: Illustration of differences in performance across demographic groups in FR with reference to different demographic factors: (a) race or ethnicity, (b) gender, and (c) age. Score distributions for mated and non-mated pairs are shown for multiple demographic groups, along with decision thresholds (dotted lines) corresponding to different operating points. This figure is intended to demonstrate that accuracy differentials may persist across a range of thresholds and are not restricted to a particular type of error (e.g., false match or false non-match).
A conventional FR system evaluates the similarity between two facial images by computing a matching score based on their extracted features to determine whether they correspond to the same individual or not. The performance of such systems is typically examined by analyzing the distributions of matching scores for image pairs of the same identity and those of different identities. Scores are thresholded into binarized decisions: match or no-match. Fig. 1 illustrates the issue of demographic bias based on race, gender, and age. Ideally, a fair FR model should exhibit equitable performance across all demographic groups (i.e., similar error rates or comparable score distributions). A commonly implicit but rarely emphasized assumption is that the groups being compared are equivalent in all aspects except for the demographic attribute of interest. For instance, comparing a group of persons wearing face masks with a group without masks- one may introduce additional variability unrelated to demographics, making the comparison unsuitable for assessing fairness. The disparity between equivalent groups contributes to the unfair or biased FR. (The formal definitions for error rates and disparities provided in Sec. II and V.) As depicted by various choices of scoring thresholds, accuracy differences exist across demographic splits and can manifest at various thresholds. This emphasizes that demographic disparities are not tied solely to false match or false non-match errors but may be present throughout the score space. This issue is further compounded by the growing reliance on FR technologies, making it imperative to identify, evaluate, and mitigate sources of accuracy differences effectively.
@@
Due to its severity and the widespread range of FR applications, demographic fairness has emerged as a crucial area of research, drawing significant attention from both the biometrics and computer vision communities [3, 4, 17, 18]. This issue has been formally incorporated into the evaluation frameworks of prominent initiatives. The National Institute of Standards and Technology (NIST) is a US agency that independently evaluates biometric technologies. Its Face Recognition Vendor Tests (FRVT) benchmark the accuracy of FR algorithms using standardized procedures and extensive datasets. Since 2019, FRVT reports have incorporated analyses of demographic disparities [19, 20], making them a key reference for assessing fairness in FR. Other initiatives, including the Maryland Test Facility (MdTF) and the European Association for Biometrics (EAB), have also contributed to this discourse, though at a smaller scale and scope compared to NIST. The MdTF, supported by the United States Department of Homeland Security (DHS), has conducted biometric technology rallies to evaluate demographic disparities in FR systems. In Europe, organizations such as the EAB have hosted dedicated events on demographic fairness in biometric systems, underlining the global importance of this topic. This research area is often positioned within the broader context of fair and trustworthy biometrics, and has been receiving substantial attention– in the form of papers, workshops, or special sessions– from leading conferences such as IEEE/CVF CVPR, WACV, IEEE FG, ICPR, ICLR, and ICML; and from reputable journals including IEEE Transactions on Information Forensics and Security (TIFS), IEEE Transactions on Biometrics, Behavior, and Identity Science (TBIOM) [21] and IEEE Signal Processing Magazine [22]. Additionally, standards organizations have recognized the need for systematic approaches. For example, the recently published ISO/IEC 19795-10 guidelines quantify demographic differentials in biometric systems, and emphasize the need for addressing fairness in FR technologies [23].
@@
Scope: Fairness and bias in machine learning are expansive topics, and their application in biometrics has drawn significant attention in recent years. Several comprehensive reviews have addressed fairness and bias in machine learning broadly [9, 24], while others focus specifically on biometrics, offering insights into various modalities such as face, fingerprint, and vein, alongside applications beyond recognition, including region of interest (ROI) detection, quality assessment, and presentation attack detection[3, 4, 17, 25]. However, as face remains the most commonly used biometric trait in operational scenarios, a substantial portion of research on differences across demographic performance or bias has concentrated on this modality.
@@
While previous reviews offer broad perspectives, the extensive literature and emerging challenges specific to demographic fairness in FR necessitate a dedicated review. In this article, we provide a consolidated discussion of recent advancements in the field, addressing the causes of demographic bias, available datasets for research, evaluation metrics for assessment of fairness, and recent mitigation techniques. Furthermore, we explore ongoing challenges that persist in addressing demographic disparities, particularly in the light of novel use cases and emerging FR applications. Although our primary focus is on race and ethnicity, as these are dominant areas of research, we also include gender-related studies within the broader context. Age-related studies, given their distinct nature (variation over time for an individual) and established body of research, are referenced only where directly relevant.
Naming Conventions: The primary objective of this review is to examine performance differences in FR across demographic groups. Here, the term performance is used broadly to encompass any metric relevant to FR evaluation, although it commonly refers to accuracy. The term differences denotes a quantifiable amount of the given performance metric measured on a given dataset. In some instances, we use the term demographic bias to describe performance differences observed across demographic groups (e.g., race, gender, age), which may be reflected in metrics such as accuracy or error rates. This terminology is used to align with common usage in recent literature and is not meant to suggest deliberate discrimination or convey any moral judgment.
@@
Additionally, terms such as race and ethnicity are often used interchangeably, although they represent distinct concepts. For clarity, we retain the terms employed by the original studies. Similarly, the names of ethnic groups vary across the literature: for instance, some works use terms like Black and White, while others prefer African (or African-American) and Caucasian. Additionally, the South Asian group is sometimes referred to as Indian, whereas Asians often refer to East Asians. To maintain consistency and respect the source material, we adhere to the original terminologies in this review. Readers are encouraged to refer to the cited works for precise definitions and context. While fairness in related topics such as face detection, image quality, expression recognition, and attribute estimation is of interest to the research community, this review exclusively focuses on demographic fairness in FR.
@@
Contributions: This review constitutes a comprehensive work dedicated to exploring demographic fairness in FR, offering a unified and holistic perspective to researchers in the field. We systematically analyze and organize key aspects, including the causes of demographic disparities, available datasets, assessment metrics, and mitigation techniques, providing a structured framework for understanding these areas. Finally, we identify emerging challenges and unresolved questions, inviting further research and innovation to advance equitability in FR systems.
@@
The structure of this paper is as follows: We begin with preliminaries of assessment of generic FR in Section II. Section III explores the causes of differences in performance, analyzing factors such as distribution of demographic groups in datasets, skin-tone, image quality, and soft attributes. We provide an overview of datasets commonly used for fairness-related research in Section IV, highlighting their demographic attributes and suitability for specific tasks. Section V reviews existing metrics for fairness evaluation, discussing their strengths and limitations. Section VI outlines recent bias mitigation strategies across different stages of the FR pipeline. We discuss open challenges and future research directions in Section VII, and conclude the review in Section VIII.
II Preliminaries of FR Evaluation
Before delving into the fairness aspect of FR systems, we briefly outline the standard FR pipeline and its associated evaluation terms. In a typical verification setting, two face images are first processed through detection and alignment stages to conform to the FR model’s input requirements. The aligned faces are then passed through a feature extractor, usually a deep neural network, which produces a compact embedding–a low-dimensional representation of the subject’s facial features. The similarity between two such embeddings is computed (commonly using cosine similarity or Euclidean distance), resulting in a continuous score typically bounded in [0, 1]. In many applications, particularly those involving enrolled users (e.g., access control), the embeddings are precomputed and stored in a secure gallery. For identification tasks, the query embedding is matched against a set of gallery templates to determine the identity.
In this context, a genuine match (or mated pair) refers to two samples originating from the same individual, while an impostor match (or non-mated pair) involves samples from different individuals. The similarity score is binarized using a threshold (often denoted as $\tau$), producing a match or non-match decision. The False Match Rate (FMR) is the proportion of impostor pairs incorrectly classified as matches (false positives), while the False Non-Match Rate (FNMR) is the proportion of genuine pairs incorrectly classified as non-matches (false negatives). These two metrics are fundamental to evaluating the accuracy and reliability of FR systems. Their values depend on the chosen threshold $\tau$, with lower FMR and FNMR generally indicating better performance. The overlap between the genuine and impostor score distributions (see Fig. 2) defines the regions where these classification errors arise.
The selection of the score threshold has a direct impact on the error rates in face recognition. Increasing the threshold tends to reduce the False Match Rate (FMR) by making it harder for impostor pairs to be incorrectly accepted as matches, but this comes at the cost of a higher False Non-Match Rate (FNMR), as more genuine pairs may be rejected, and vice versa. It is important, however, to distinguish between the inherent overlap in the score distributions and the binary classification decision that results from thresholding. The overlap in the range of similarity scores between mated and non-mated pairs indicates intrinsic ambiguity in the matching process, while the threshold simply determines the point of separation between match and non-match decisions.
Two demographic groups may exhibit a similar degree of score overlap yet yield different FMR and FNMR values due to differences in the shapes of their score distributions or disparities in group-specific population characteristics. Conversely, distributions across groups may differ significantly, while their overall error rates remain similar. These variations– in either score distribution overlap or resulting error rates– across demographic subgroups form the foundation of what is commonly referred to as bias in FR systems. This review focuses on analyzing, evaluating, and mitigating these disparities.
III Causes of Performance Differences
@@
In this section, we consolidate findings from existing works related to causes of performance differences in FR. Considering the wide range of research in this area, we have grouped the causes into categories for clarity and systematic review. These categories encompass factors such as imbalances in training datasets, variability in skin-tones, algorithmic sensitivity, image quality, related covariates, combined or intersectional factors, and soft attributes. Though this categorization simplifies the organization, it is important to note that many studies attribute accuracy differences to multiple, overlapping factors, making strict classification difficult. We have categorized works based on their primary focus or findings, but some studies have been referenced across multiple categories to reflect their broader relevance. This approach also highlights that causes of demographic differences or bias are inherently multifaceted and interconnected, requiring a thorough understanding to analyze and mitigate these disparities. A quick summary of various works discussed in this section can be found in Table I.
III-A Training Datasets
@@
Early studies in this area speculated that the issue of performance differences in FR systems across demographic groups often stems from imbalanced or unrepresentative datasets, significantly influencing both training and evaluation outcomes. Research by Krishnapriya et al. [26] demonstrated how demographic groups, such as African-American cohorts, exhibit higher FMRs, while Caucasian cohorts face higher FNMR, highlighting the interplay between race and matching thresholds. The FRVT conducted by NIST [19] reported increased false positive identifications in women, children, and the elderly, alongside higher false negatives in under-represented racial groups– particularly Black, East Asian, and Native American individuals– emphasizing the intricate interactions between dataset characteristics and demographic attributes.
@@
In a pre-deep learning era, studies like Klare et al.  [27] examined the impact of demographic balance in training datasets on FR performance. While their results indicated that underrepresented groups– such as females and Black individuals– benefited from increased representation, they also found that balancing the dataset alone did not fully eliminate disparities in recognition accuracy at test time. To further mitigate this, they proposed training exclusive cohort-specific models and using dynamic model selection during inference, highlighting that data balance might often be complemented with architectural or procedural adjustments to meaningfully reduce accuracy differences. Cavazos et al. [28] identified how dataset complexity and identification thresholds contribute to racial bias, such as the need for higher thresholds for East-Asian faces to achieve comparable false acceptance rates (FARs). Gwilliam et al. [29] challenged the prevailing assumptions about the necessity of balanced datasets by demonstrating that skewed distributions favoring African faces reduced racial differentials more effectively than balanced datasets. In [30, 31], Albiero et al. demonstrated that balancing training data for both genders (male and female) did not help reducing the corresponding differential in FR accuracy. Wu and Bowyer [32] expanded this discussion, emphasizing that mere balance in identities or number of images is insufficient to address disparities, highlighting additional factors like brightness and head pose during dataset assembly. These factors–often referred to as soft attributes–include physical traits, behavioral patterns, or material accessories associated with a person. They are descriptive and can be categorized or classified [33]. The soft-biometric attributes tend to influence the performance of FR systems.
@@
Other works delved into specific aspects of demographic balance in datasets. Wang et al.  [34] observed that even race-balanced datasets failed to eliminate accuracy differentials, hypothesizing that certain ethnicities are inherently more challenging to recognize. Kolla and Savadamuthu [35] highlighted the influence of facial quality and racial feature gradations on model fairness. Focusing on inter-sectional bias, Muthukumar et al.  [36] identified structural facial features as significant contributors, particularly for dark-skinned females, over attributes like skin-tone or hair length. Cook et al. [37] further analyzed the role of image acquisition conditions, noting how factors such as skin reflectance and environmental conditions disproportionately affect darker-skinned individuals, thus advocating for standardized acquisition protocols to mitigate differences in FR performance.
@@
III-B Variability in Skin-tone
The influence of skin-tone on the performance of FR systems has been extensively studied, revealing significant demographic disparities. A common method to categorize skin-tone for FR studies is the Fitzpatrick Skin Type Classification, which divides skin types into 6 categories (Type I to VI) based on skin’s reaction to ultraviolet (UV) exposure and general pigmentation [38]. While widely used due to its simplicity, it was originally designed for dermatology, not computer vision. Alternative categorization schemes include the von Luschan’s chromatic scale, a 36-point system based on observed skin color using standardized tiles, and the Monk Skin-Tone (MST) Scale, a 10-tone scale designed to represent a broader and more inclusive range of skin-tones [39]. Each scheme varies in granularity and applicability, and ongoing research continues to assess their effectiveness.